OpenStudy (anonymous):

Nancy can you help me Radium decreases at the rate of 0.0428 percent per year. a. What is its half-life? (A half-life of a radioactive substance is defined to be the time needed for half of the material to dissipate. b. Write a recurrence relation to describe the decay of radium, where rn is the amount of radium remaining after n years.

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

y = Ae^(-kt)

OpenStudy (anonymous):

general formula for exponential decay

OpenStudy (anonymous):

A/2=A(0.000428)^x 1/2=(0.000428)^x find x

OpenStudy (anonymous):

* correction A/2=A(1-0.000428)^x

OpenStudy (anonymous):

\[1/2=(1-0.000428)^x\]

OpenStudy (anonymous):

\[\log _{0.999572}(.5)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that's half life

OpenStudy (anonymous):

so how does translate to what I need for b

OpenStudy (anonymous):

\[r_n=r_{n-1}(0.999572)\]

OpenStudy (anonymous):

are you electrical engineer or student?

OpenStudy (anonymous):

student this is module for 1 credit and I'm struggling!

OpenStudy (anonymous):

I was asking elecengineer

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

np

OpenStudy (anonymous):

by the way, did you figure out that credit card problem?

OpenStudy (anonymous):

Suppose that one started out with 2 grams of radium. Find a solution for the discrete dynamical system illustrating this process and give the value for r(100), the amount remaining after 100 years.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, we can use above formula =2(1-0.000428)^100

OpenStudy (anonymous):

approximately 1.92

OpenStudy (anonymous):

that make sense because its half life in 1619 years so in 100 years there is not much difference